# VisionUnite: A Vision-Language Foundation Model for Ophthalmology Enhanced with Clinical Knowledge

### v1

@@ Abstract
The need for improved diagnostic methods in ophthalmology is acute, especially in the less developed regions with limited access to specialists and advanced equipment. Therefore, we introduce VisionUnite, a novel vision-language foundation model for ophthalmology enhanced with clinical knowledge. VisionUnite has been pretrained on an extensive dataset comprising 1.24 million image-text pairs, and further refined using our proposed MMFundus dataset, which includes 296,379 high-quality fundus image-text pairs and 889,137 simulated doctor-patient dialogue instances. Our experiments indicate that VisionUnite outperforms existing generative foundation models such as GPT-4V and Gemini Pro. It also demonstrates diagnostic capabilities comparable to junior ophthalmologists. VisionUnite performs well in various clinical scenarios including open-ended multi-disease diagnosis, clinical explanation, and patient interaction, making it a highly versatile tool for initial ophthalmic disease screening. VisionUnite can also serve as an educational aid for junior ophthalmologists, accelerating their acquisition of knowledge regarding both common and rare ophthalmic conditions. VisionUnite represents a significant advancement in ophthalmology, with broad implications for diagnostics, medical education, and understanding of disease mechanisms.
Introduction
In recent years, the global prevalence of ophthalmic diseases has surged beyond 2.2 billion, with over 1 billion individuals experiencing visual impairment due to limited access to essential medical services for conditions like myopia, hyperopia, glaucoma, and cataracts. The critical situation primarily stems from a shortage of ophthalmologists in low-income and middle-income regions, resulting in inadequate provision of ophthalmic services. Compounding this challenge, the World Health Organization’s World Vision Report estimates a staggering cost of 14.3 billion to address these issues, underscoring the financial burden. Consequently, there is an escalating need for swift and precise comprehensive diagnoses facilitated by existing artificial intelligence technology. Notably, numerous researchers have made strides in developing vision models for diagnosing eye diseases, exemplified by works such as the works [1, 2, 3, 4, 5, 6]. The development of deep learning models for ophthalmology faces three significant challenges. Firstly, these models are often tailored to diagnose specific diseases and lack the ability to provide comprehensive assessments for multiple conditions simultaneously. It is a critical shortfall, as patients frequently suffer from multiple ailments at once, particularly in older populations. For instance, as reported by the American Academy of Ophthalmology [7], it is not uncommon for individuals aged 65 and above to have more than one eye disease. Secondly, there is a persistent issue with effective user interaction, which is essential for practical clinical implementation. Thirdly, many of these AI models lack interpretability in their diagnostic results, which is crucial for trust and reliability in medical settings. In an ideal scenario, a large vision-language model would seamlessly navigate diverse clinical scenarios, including disease screening, diagnostic process optimization, and junior ophthalmologist training. Such a model would integrate visual and linguistic data effectively, aligning closely with the diagnostic criteria used by medical professionals and adhering to clinical consensus guidelines. This approach would ideally involve identifying the lesion area and type before proceeding with a diagnosis, enhancing both the interpretability and accuracy of medical evaluation. However, prevailing models primarily fall short of this ideal. Current vision models such as RETFound [8], typically only offer diagnostic predictions by identifying the types of diseases without providing the necessary explanatory context. It results in a lack of interpretability that diverges from the standards of medical practice. Similarly, while there are advanced language models referenced in the study [9], there remains a significant gap in effective vision-language interaction. This deficiency is also noted in the lack of comprehensive user interaction which hinders the model from further understanding the needs of users, as described in work [10]. These limitations impair the accuracy and universality of diagnoses made by models, underscoring the need for an integrated approach that can effectively combine visual and linguistic information.
@@ Introduction
To meet the above requirements, we introduce VisionUnite, a Large Vision Language Model tailored for Ophthalmology with Clinical Knowledge. As depicted in Figure 1 (a), VisionUnite addresses three critical challenges: 1) the inability to predict open-ended multiple diseases or conditions, 2) the lack of effective user interaction, and 3) diagnostic results with limited interpretability. Regarding the first challenge, VisionUnite demonstrates the capability to predict a wide range of lesion types in fundus images during the initial responses, encompassing conditions like Retinal Hemorrhage and Macular Edema. It stands in contrast to existing models constrained by data size and disease types, precluding open-ended disease prediction. Addressing the challenge of the absence of user interaction, VisionUnite implements multiple rounds of dialogue and effectively follows user instructions. The multiple rounds of dialogue enhance the comprehension of models to user commands and image information and contribute to the generation of more robust diagnostic reports. Adopting the multi-round dialogue can significantly bolster the interpretability of VisionUnite’s diagnoses. The responses of VisionUnite offer detailed explanations for diagnoses (retinal hemorrhage and macular edema) and propose future treatment plans.
Results

### v2

@@ Abstract
The need for improved diagnostic methods in ophthalmology is acute, especially in the underdeveloped regions with limited access to specialists and advanced equipment. Therefore, we introduce VisionUnite, a novel vision-language foundation model for ophthalmology enhanced with clinical knowledge. VisionUnite has been pretrained on an extensive dataset comprising 1.24 million image-text pairs, and further refined using our proposed MMFundus dataset, which includes 296,379 high-quality fundus image-text pairs and 889,137 simulated doctor-patient dialogue instances. Our experiments indicate that VisionUnite outperforms existing generative foundation models such as GPT-4V and Gemini Pro. It also demonstrates diagnostic capabilities comparable to junior ophthalmologists. VisionUnite performs well in various clinical scenarios including open-ended multi-disease diagnosis, clinical explanation, and patient interaction, making it a highly versatile tool for initial ophthalmic disease screening. VisionUnite can also serve as an educational aid for junior ophthalmologists, accelerating their acquisition of knowledge regarding both common and underrepresented ophthalmic conditions. VisionUnite represents a significant advancement in ophthalmology, with broad implications for diagnostics, medical education, and understanding of disease mechanisms. The source code is at https://github.com/HUANGLIZI/VisionUnite.
Index Terms:
Foundation Model, Generative AI, Multimodal
I Introduction
In recent years, the global prevalence of ophthalmic diseases has surged beyond 2.2 billion, with over 1 billion individuals experiencing visual impairment due to limited access to essential medical services for conditions like myopia, hyperopia, glaucoma, and cataracts. The critical situation primarily stems from a shortage of ophthalmologists in low-income and middle-income regions, resulting in inadequate provision of ophthalmic services. Compounding this challenge, the World Health Organization’s World Vision Report estimates a staggering cost of 14.3 billion to address these issues, underscoring the financial burden. Consequently, there is an escalating need for swift and precise comprehensive diagnoses facilitated by existing artificial intelligence (AI) technology.
Notably, numerous researchers have made strides in developing vision models for diagnosing eye diseases, exemplified by works such as the works [1, 2]. Current AI-based models for ophthalmology face three significant challenges: disease-specific diagnosis limitations, ineffective user interactions, and lack of result interpretability. Firstly, these models are often tailored to diagnose specific diseases and cannot provide comprehensive assessments for multiple conditions simultaneously. It is a critical shortfall, as patients frequently suffer from multiple ailments, particularly in older populations. For instance, as reported by the American Academy of Ophthalmology [3], it is not uncommon for individuals aged 65 and above to have more than one eye disease. Secondly, there is a persistent issue with effective user interaction, which is essential for practical clinical implementation. Thirdly, many of these AI models lack interpretability in their diagnostic results, which is crucial for trust and reliability in medical settings.
An ideal solution would involve a comprehensive large vision-language model seamlessly managing diverse clinical scenarios, including disease screening, diagnostic process optimization, and junior ophthalmologist training. Such a model would integrate visual and linguistic data effectively, aligning closely with the diagnostic criteria used by medical professionals and adhering to clinical consensus guidelines. This approach would ideally involve identifying the lesion area and type before proceeding with a diagnosis, enhancing both the interpretability and accuracy of medical evaluation. However, prevailing models primarily fall short of this ideal. Current vision models like RETFound [4] often diagnose diseases without explaining their findings, which lacks interpretability and deviates from medical standards. Advanced language models show a significant shortfall in effective vision-language integration [5, 6]. This issue extends to poor user interaction and limits model responsiveness to user needs [7]. These limitations impair the accuracy and universality of diagnoses made by models, underscoring the need for an integrated approach that can combine visual and linguistic information.
@@ I Introduction
We introduce VisionUnite, a large vision language model tailored for ophthalmology, incorporating extensive clinical knowledge to effectively address these challenges. As depicted in Figure 1 (a), VisionUnite addresses three critical challenges: 1) the inability to predict open-ended multiple Diseases, 2) the lack of effective user interaction, and 3) diagnostic results with limited interpretability. Regarding the first challenge, VisionUnite demonstrates the capability to predict a wide range of lesion types in fundus images during the initial responses, encompassing conditions like Retinal Hemorrhage and Macular Edema. It stands in contrast to existing models constrained by data size and disease types, precluding open-ended disease prediction. Addressing the challenge of the absence of user interaction, VisionUnite implements multiple rounds of dialogue and effectively follows user instructions. The multiple rounds of dialogue enhance the comprehension of models to user commands and image information and contribute to the generation of more robust diagnostic reports. Finally, VisionUnite improves interpretability by providing detailed explanations of diagnoses, as illustrated in Figure 1. These detailed responses help clinicians better understand the rationale behind the AI’s decisions, thereby fostering greater trust in the model’s diagnostic outputs.
Figure 1: (a) Previous vision models could only diagnose specific diseases as positive or negative, lacking the ability to provide clinical explanations or interact with patients. In contrast, our proposed VisionUnite represents a new paradigm. It can predict a wide range of diseases and allows real-time conversations with users, incorporating their feedback. Furthermore, VisionUnite offers clinically relevant explanations in its output, making it more understandable and useful. (b) The label distribution of the proposed MMFundus dataset, which includes eight main categories excluding the ”Others” class.
Figure 2: (a) VisionUnite is built with a transformer-based vision encoder and a specialized vision adapter designed for classifying six different signs including Vascular, Macular, FBC (Fundus Boundary Color), OCD (Optical Cup Disc), FHE (Fundus Hemorrhages Examination), and Others. It includes a vision projector to align visual embeddings with text tokens. (b) The illustration of image-text contrastive learning (CLIP Loss). (c) The illustration of classification supervised learning (CLS Loss). (d) The illustration of text-generation supervised learning (LLM Loss).
II Related Work
Vision-Language Foundation Models (VLFMs) have become a cornerstone in multimodal AI, skillfully blending vision and language domains to comprehend and generate multimodal content. These models, as discussed in references such as [8, 9], are trained on extensive datasets featuring images paired with textual annotations, facilitating high proficiency in tasks like image captioning, visual question answering, and cross-modal retrieval. The use of transformer architectures is central to their ability, allowing them to adeptly manage the intricacies of multimodal inputs [10, 11]. Recent innovations like the LLaMA-Adapter[10] push these capabilities further by introducing adaptability in processing diverse datasets, enhancing the model’s flexibility and application range. Similarly, LLaVa-Med[12] extends these principles into the medical domain, integrating varied data sources to construct a comprehensive view of patient health, thus aiding better clinical decision-making. Despite these advancements, significant challenges persist. The complexity inherent in VLFMs often results in a ”black box” scenario, particularly problematic in healthcare settings where interpretability is crucial  [5, 7]. Models like Lama-adapter and Llava-med, while innovative, do not fully address the lack of transparency in how decisions are derived, which can be a barrier to their trust and utility in clinical practice. InternVL [13] represents a novel approach to generic visual-linguistic tasks. However, the complexity of the multi-stage training process, especially with layers of contrastive and generative training, can make it challenging to trace the specific decision process. Therefore, we propose the vision adapter to align coarse-grained labels explicitly. Additionally, there remains a significant gap in effective vision-language interaction. The current models still struggle to mimic the human interaction between visual cues and linguistic context, which limits the dynamic interactions.
III Method
III-A Overview